For very large resistances it is easy to construct R-C circuits that have time constants of several seconds or minutes. How might this fact be used to measure very large resistances, those that are too large to measure by more conventional means?


1
Expert's answer
2022-02-28T10:42:16-0500

Answer

using time constant formula

"\tau=RC=\frac{1}{2\pi f}"

So resistance is given by

"\R=\frac{1}{2\pi fC}"